# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

### v3

@@ Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. Such equivariance constraint often requires specialized architecture designs to achieve effective predictions, inevitably introducing tremendous computational costs. Canonicalization, a classical technique for geometry, has recently been explored for efficient learning with symmetry. In this work, we revisit the problem of crystal tensor property prediction through the lens of canonicalization. Specifically, we demonstrate how polar decomposition, a simple yet efficient algebraic method, can serve as a form of canonicalization and be leveraged to ensure equivariant tensor property prediction. Building upon this insight, we propose a general $O(3)$-equivariant framework for fast and accurate crystal tensor property prediction, referred to as GoeCTP. By utilizing canonicalization, GoeCTP achieves high efficiency without requiring the explicit incorporation of equivariance constraints into the network architecture. Experimental results indicate that GoeCTP achieves the best prediction accuracy and runs up to 13$\times$ faster compared to existing state-of-the-art methods in benchmarking datasets, underscoring its effectiveness and efficiency.
1 Introduction
@@
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering [51]. These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders. Accurate prediction of these tensor properties is critical for novel materials discovery and design with targeted characteristics. Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) [35]. These classical simulation techniques can provide acceptable error margin for predicting various material properties. However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons [51], hindering their applicability in practice.
@@
Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) [1, 26, 2, 47] or transformers [50, 42, 49, 20, 44]. Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures. However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress [33, 38, 51]; their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance [51].
@@
Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures [28, 25, 11, 46, 34, 51, 53]. These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds [22], n-body simulation [17], and antibodies generation [29], is the use of “canonicalization". Specifically, canonicalization maps a geometric data to an invariant representation [27, 6], referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the network. However, to date, it has not been explored for crystal tensor property prediction.
In this work, we revisit the task of crystal tensor property prediction through the lens of canonicalization and introduce a novel canonicalization strategy tailored for this particular setting. Specifically, rather than embedding equivariance directly into model architecture, we propose a simple yet effective canonicalization module, termed R&R. In particular, our canonicalization is instantiated based on polar decomposition, a continuous mapping technique [6] that can provide enhanced robustness. During prediction, the R&R module applies rotations and reflections to transform the input crystal structure into its canonical form. The canonical form is then fed into a property prediction network to obtain the corresponding canonical tensor representation. Simultaneously, the orthogonal matrix derived from the R&R module is utilized to recover the equivariant output via the tensor transformation rule, enabling equivariant tensor property prediction without incurring additional computational overhead. We conducted experiments on dielectric, piezoelectric, and elastic tensor datasets, respectively, showcasing that the proposed method can achieve $O(3)$-equivariant tensor prediction while maintaining high efficiency. Compared to the previous state-of-the-art work Yan et al. [51], the GoeCTP method achieves higher prediction accuracy and runs by up to 13$\times$ faster in benchmarking datasets.
2 Background
2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters [49, 45, 16]. General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
@@ 2.1 Preliminaries
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA [30], Yan et al. [51], and VELAS [37].
2.2 Problem statement
@@
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature [28, 46, 51], our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$ group. The $O(3)$ group is an orthogonal group, consisting of rotations and reflections. The elements $g\in O(3)$ act on vectors or tensors in $\mathcal{V}$, $\mathcal{W}$ through their respective group representations, $\rho_{\mathcal{V}}:O(3)\to GL(\mathcal{V})$ and $\rho_{\mathcal{W}}:O(3)\to GL(\mathcal{W})$, where $GL(\mathcal{V})$ and $GL(\mathcal{W})$ are the space of invertible linear maps $\mathcal{V}\to\mathcal{V}$ and $\mathcal{W}\to\mathcal{W}$, respectively. Specifically, the action of an element $g\in O(3)$ on crystal data $\mathbf{M}$ is defined as: $g\cdot\mathbf{M}=\rho_{\mathcal{V}}(g)\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})$. For dielectric tensor property $\mathcal{Y}$, the action of $g$ is given by: $g\cdot\mathcal{Y}=\rho_{\mathcal{W}}(g)\mathcal{Y}=\mathbf{Q}\mathcal{Y}\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is a $3\times 3$ orthogonal matrix (for additional details regarding the transformation of tensor properties, please refer to Appendix A.8).
$\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction. In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies [13, 48, 52, 40]. Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
@@
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix [51]. For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1 (for more equivariance details, see Appendix A.8).
3 Methodology
@@
3.1 The Motivation of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work gather the rotational information from the neighboring nodes to the central node $i$ through the following computation formulated [51]:
$\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),$ (3)
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ [7], and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. To achieve $O(3)$ equivariance of the model and ensure high prediction accuracy for high-order tensor prediction tasks, such as elastic tensor prediction, the rotation order $\lambda$ is set to 3 [51], i.e., $\mathbf{Y}_{3}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{7}$. In contrast, for lower-order tensor prediction tasks, setting $\lambda$ to 1 is sufficient to achieve the model’s $O(3)$ equivariance, i.e., $\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{3}$. Compared to lower-order tensor prediction, the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ nearly doubles for high-order tensor predictions, leading to a significant increase in computational complexity. Consequently, this model architecture introduces tremendous computational costs for tensor property predictions. If we can achieve the model’s $O(3)$ equivariance and maintain the model performance for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, or even without relying on harmonic decomposition, computational costs could be drastically reduced. Inspired by the recent exploration of using canonicalization in efficient learning with symmetry, we aim to propose a principled approach based on canonicalization to enable $O(3)$-equivariant tensor predictions for crystal materials.
3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques [27, 17, 6].
Definition 3.1.
(Orbit.) The orbit of a crystal $\mathbf{M}$ is defined as $\text{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
@@ 3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
Definition 3.2.
(Orbit Canonicalization.) A function denoted as $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization, if $\forall\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$ and $\exists\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M}),$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$, where $\mathbf{M}_{0}$ is the canonical form of the orbit.
@@
Although an $O(3)$-invariant prediction function has been established via orbit canonicalization, the complete process of equivariant prediction requires specific group transformations that map all crystal configurations within the orbit to their canonical form. To address this requirement, we introduce the concept of rigid registration. Rigid registration is a fundamental concept in fields such as computer vision and computational geometry [41]. It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object. This concept can be naturally adapted to the crystal, as follows.
Definition 3.3.
@@
Using Definition 3.3, we can identify a transformation that matches two configurations of a crystal. Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form. By leveraging Definition 3.2 and Definition 3.3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
Proposition 3.4.
($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function formulated as $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
The proof of Proposition 3.4 is provided in Appendix A.6. Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3.3, to find specific group transformations that map crystal configurations within the orbit to their canonical form. This canonical form is determined through orbit canonicalization, as described in Definition 3.2. Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance. That is, under this perspective, there are no specific requirements or special treatments needed for the network design, as the $O(3)$ transformations applied to the crystal can be externally determined and simultaneously applied directly to the output. Building upon Proposition 3.4, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
3.3 Our Proposed Framework: GoeCTP
@@
Theorem 3.5.
(Polar Decomposition [9, 12, 16].) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a Hermitian positive semi-definite matrix.
In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form. Therefore, throughout this work, we adopts the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Then building on Theorem 3.5 , we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
Proposition 3.6.
@@
The detailed proof of Proposition 3.6 is provided in Appendix A.7. Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$. Thus, $\mathbf{H}$ can serve as the canonical form. Moreover, the rationale of using polar decomposition is that it is a continuous canonicalization, which can provide improved robustness [6]. For clarity, we provide details on this continuity in Appendix A.14, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.15, respectively. Building upon Proposition 3.6, we utilize polar decomposition to construct the proposed GoeCTP model.
R&R Module. The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation. Therefore, based on Proposition 3.6, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization. Formally, the R&R Module can be expressed as:

### v4

@@ Abstract
[…]
Introduction
@@
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (yanspace). These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders. Accurate prediction of these tensor properties is critical for novel materials discovery and design with targeted characteristics. Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) (petousis2016benchmarking). These classical simulation techniques can provide acceptable error margin for predicting various material properties. However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons (yanspace), hindering their applicability in practice.
@@
Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (chen2019graph; louis2020graph; choudhary2021atomistic; xie2018crystal) or transformers (yancomplete; taniaicrystalformer; yan2022periodic; lee2024density; wang2024conformal; ito2025rethinking). Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures. However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (nye1985physical; resta1994macroscopic; yanspace); their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance (yanspace).
@@
Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures (mao2024dielectric; lou2024discovery; heilman2024equivariant; wen2024equivariant; pakornchote2023straintensornet; yanspace; zhong2023general). These methods generally integrate directional features and spherical harmonics to preserve equivariance. However, achieving this often requires computationally expensive operations such as tensor products, which introduce substantial overhead, especially when dealing with high-order tensor data. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
@@
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. The visualization of the crystal structures on the left was generated using VESTA (momma2011vesta), while the visualization of the crystal tensor properties on the right follows the method of yanspace and VELAS (ran2023velas).
One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds (linequivariance), n-body simulation (kaba2023equivariance), and antibodies generation (martinkus2023abdiffuser), is the use of “canonicalization”. Specifically, canonicalization maps a geometric data to an invariant representation (ma2024canonicalization; dymequivariant), referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the backbone network. However, to date, it has not been explored for crystal tensor property prediction.
[…]
Background
Preliminaries
Crystalline materials consist of a periodic arrangement of atoms in 3D space, characterized by a repeating unit known as the unit cell. A complete crystal can thus be described by the atomic types, atomic coordinates, and lattice parameters of a single unit cell (yan2022periodic; wang2024crystalline; jiaospace; huangcode; hua2025local). Mathematically, a crystal can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes the atomic features of $n$ atoms, $\mathbf{X}=[\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains their Cartesian coordinates, and $\mathbf{L}=[\boldsymbol{l}_{1},\boldsymbol{l}_{2},\boldsymbol{l}_{3}]\in\mathbb{R}^{3\times 3}$ is the lattice matrix composed of three basis vectors. The entire crystal can then be expressed as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\boldsymbol{a}_{i}},\hat{\boldsymbol{x}}_{i})|\hat{\boldsymbol{x}}_{i}=\boldsymbol{x}_{i}+k_{1}\boldsymbol{l}_{1}+k_{2}\boldsymbol{l}_{2}+k_{3}\boldsymbol{l}_{3},\hat{\boldsymbol{a}_{i}}=\boldsymbol{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$, which enumerates all atomic positions generated by lattice translations.
Alternatively, using the lattice matrix $\mathbf{L}$ as the basis vectors leads to the fractional coordinate representation, where each atom is assigned a fractional coordinate $\boldsymbol{f_{i}}=[f_{i,1},f_{i,2},f_{i,3}]^{\top}\in[0,1)^{3}$ and its Cartesian coordinate is given by $\boldsymbol{x}_{i}=\sum_{j}f_{i,j}\boldsymbol{l}_{j}$. Accordingly, the crystal can also be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\boldsymbol{f}_{1},\cdots,\boldsymbol{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ denotes the fractional coordinates of all atoms in the unit cell.
Problem statement
@@
[…]
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As previously proposed in the literature (mao2024dielectric; wen2024equivariant; yanspace), our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$ group.
The $O(3)$ group is an orthogonal group, consisting of rotations and reflections. The elements $g\in O(3)$ act on vectors or tensors in $\mathcal{V}$, $\mathcal{W}$ through their respective group representations, $\rho_{\mathcal{V}}:O(3)\to GL(\mathcal{V})$ and $\rho_{\mathcal{W}}:O(3)\to GL(\mathcal{W})$, where $GL(\mathcal{V})$ and $GL(\mathcal{W})$ are the space of invertible linear maps $\mathcal{V}\to\mathcal{V}$ and $\mathcal{W}\to\mathcal{W}$, respectively. Specifically, the action of an element $g\in O(3)$ on crystal data $\mathbf{M}$ is defined as: $g\cdot\mathbf{M}=\rho_{\mathcal{V}}(g)\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})$. For dielectric tensor property $\mathcal{Y}$, the action of $g$ is given by: $g\cdot\mathcal{Y}=\rho_{\mathcal{W}}(g)\mathcal{Y}=\mathbf{Q}\mathcal{Y}\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is a $3\times 3$ orthogonal matrix (for additional details regarding the transformation of tensor properties, please refer to Appendix A.4).
$\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction.
In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (hoogeboom2022equivariant; xuequivariant; zheng2024relaxing; song2024equivariant; aykent2025gotennet; cen2024high). Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq.1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
@@
where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix (yanspace). For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig.1 (for more equivariance details, see Appendix A.4).
Methodology
@@
The Motivation of Our Framework
To enforce $O(3)$ equivariance as defined in Eq. 2, existing crystal tensor prediction approaches typically rely on irreducible representations and Clebsch-Gordan (CG) tensor products. Although these formulations are theoretically rigorous and have shown strong empirical performance, they introduce substantial computational overhead due to the complexity of tensor operations.
Recently, several studies have explored the scalarization trick to exploit high-degree steerable features, thereby eliminating the need for CG tensor products while maintaining both accuracy and efficiency (aykent2025gotennet; cen2024high; han2025survey). Similarly, if $O(3)$ tensor equivariance could be achieved for high-order tensor predictions without relying on tensor product operations, the computational cost could be significantly reduced. Motivated by recent progress in canonicalization-based symmetric learning, we propose a principled framework that leverages canonicalization to realize $O(3)$-equivariant tensor prediction for crystalline materials.
$O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques (ma2024canonicalization; kaba2023equivariance; dymequivariant).
Definition 1 (Orbit.).
The orbit of a crystal $\mathbf{M}$ is defined as $\mathrm{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
@@
Definition 2.
(Orbit Canonicalization [^1] For simplicity, we refer to the procedure as canonicalization, though it may more precisely be regarded as a quasi-canonicalization that relaxes permutation invariance. ) A function $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization if there exists a canonical form $\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M})$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$ for all $\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$.
@@
Although an $O(3)$-invariant prediction function has been established via orbit canonicalization, the complete process of equivariant prediction requires specific group transformations that map all crystal configurations within the orbit to their canonical form. To address this requirement, we introduce the concept of rigid registration. Rigid registration is a fundamental concept in fields such as computer vision and computational geometry (tam2012registration). It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object. This concept can be naturally adapted to the crystal, as follows.
Definition 3.
@@
Using Definition 3, we can identify a transformation that matches two configurations of a crystal. Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form. By leveraging Definition 2 and Definition 3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
Proposition 1.
($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we define a new function $h(\mathbf{M})=R_{\mathbf{M}}\!\left(\mathbf{M},C_{\mathbf{M}}(\mathbf{M})\right)\!\cdot\!f\!\left(C_{\mathbf{M}}(\mathbf{M})\right)$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
The proof of Proposition 1 is provided in Appendix A.3. Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3, to find specific group transformations that map crystal configurations within the orbit to their canonical form. This canonical form is determined through orbit canonicalization, as described in Definition 2. Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance. Under this perspective, our framework introduces no architectural constraints or special design requirements on the backbone network itself, since the $O(3)$ transformations of the crystal are externally determined through canonicalization and applied directly to the outputs. Building upon Proposition 1, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
Our Proposed Framework: GoeCTP
@@
Theorem 1.
(Polar Decomposition (hall2013lie; higham1986computing; jiaospace).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a symmetric positive semi-definite matrix.
In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form. Therefore, throughout this work, we adopt the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Then building on Theorem 1, we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
Proposition 2.
@@
The detailed proof of Proposition 2 is provided in Appendix A.3. Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$. Thus, $\mathbf{H}$ can serve as the canonical form. Moreover, using polar decomposition is a continuous canonicalization, which can provide improved robustness (dymequivariant). For clarity, we provide details on this continuity in Appendix A.9, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.10, respectively. Building upon Proposition 2, we utilize polar decomposition to construct the proposed GoeCTP model.
R&R Module.
The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation. Therefore, based on Proposition 2, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization. Formally, the R&R Module can be expressed as: